# Simulation of Quantum Computers: Review and Acceleration Opportunities

### v1

@@ 1. Introduction
While several simulators have recently been deployed and made publicly available, the need to simulate more complex algorithms, circuit configurations, or more accurately model qubit behavior leads to an exponential increase in computational and memory demand for the simulations. Scaling up the simulations without increasing exponentially the amount of necessary resources, requires optimisations or approximations at different levels. As such, several classical computer techniques, such as data compression or optimized parallel execution, have been applied to reduce the memory requirements and accelerate the computations (Lykov et al., 2022; Bayraktar et al., 2023).
@@
The organization of this survey is as follows. Section 2 introduces the essential concepts in quantum computing necessary for understanding the rest of the sections. In Section 3 we give an overview of the various parts of a quantum computer. Section 4 covers simulations at different levels and provides an overview of the available simulators for each of them, also introducing the main bottlenecks for scaling up the number of simulated qubits. Section 5 details the most common hardware platforms used for simulation. The following sections cover different acceleration methods, sorted by platform: CPU in Section 6, GPU in Section 7, and FPGA in Section 8. A summary and indication of future directions for hardware-aware optimization is presented in Section 9 and Section 10 concludes this work.
@@ 2.2.1. Single-qubit gates
These matrices correspond to the rotation of $\pi$ radiant around respectively the x,y, and z axis of the Bloch sphere.
@@ 2.3. Quantum circuit execution
[Two qubit circuit, similar to two single-qubit circuit stacked vertically.]The image contains two quantum circuit diagrams, each depicting a different quantum gate operation applied to qubit states. The first diagram shows an initial quantum state psi on the left side of the X gate, represented as a rectangular box. The output of this operation is a transformed quantum state psi first on the right side of the box. The second diagram is similar, showing a quantum state labelled phi connected to the quantum gate Y also represented as a box. The output is a transformed quantum state labelled phi’.
@@ 3. Quantum computers
[Quantum computer control electronics(host, pulse generation, readout), interface blocks (DAC, ADC) and quantum chip]The image shows how the different parts of the system are connected. The different parts are placed in a elliptic configuration, starting from the host on the left and proceeding counter-clockwise back to the host. Starting from the host, represented as a computer symbol. An arrow goes from the host to the next component, the pulse generation block which is represented as a rectangle box, situated on the bottom right of the host. The pulse generation block is then connected via an arrow to the base of the triangle used to represent the digital-to-analog converter. The point of the triangle, on the opposite side of the base, is on the right side and it is connected with an arrow to the quantum chip, which is on the rightmost side of the figure. The quantum chip, represented as a chip symbol, is then connected to the analog-to-digital converter, which is represented as the same triangle as the DAC, but pointing in the opposite direction, and it is placed on the top left of the quantum chip. The ADC is then connected to the readout, represented as a rectangle box, and it is then connected back to the host.
@@ 3.1. Quantum chip
As already mentioned in the introduction, one of the main limitations of current physical qubits is their short coherence time. This is the timescale of the exponential decay of the qubit superposition state (Abad et al., 2022). The correctness of quantum computers is usually measured in terms of a fidelity for the operations or the output state.
@@ 4. Simulations of quantum computers
  • Algorithmic level — this is the highest level of the system, where algorithms are mappped to quantum circuits. This is comparable to the functional-level classical simulation.
@@ 4.2. Gate level
[The figure represents two coupled qubits, each depicted as a Josephson junction and a capacitance, and connected between them by a capacitor]Each qubit is shown as a block with the labels omega1 or omega2 and alpha1 or alpha2. Each block has a Jhosepson junction on the left, represented as a square box with lines crossing from corner to corner, and a capacitor on the right, represented with the standard symbol. On the bottom the box is shown as connected to ground and on the top is connected to the other box through a capacitive coupling, shown as a capacitor labeled g12. The qubits are placed one on the left and one on the right, with the capacitive coupling in the top middle of the figure.
@@ 4.3. Algorithmic level
  • Intel Quantum SDK (Khalate et al., 2022): Provides two different simulators with two different modes of simulation, Generic Qubits and Intel Hardware. The Generic Qubits simulation is a state-vector-based simulation, which uses as a backend the Intel Quantum Simulator, allowing for a qubit-agnostic execution. The Intel Hardware mode uses as its core the Quantum Dot Simulator, which simulates the Intel quantum hardware, currently under development.
@@ 8. Acceleration using FPGA
Although FPGAs are more limited in terms of resources and frequency when compared to CPUs and GPUs, they allow for more flexibility, leading to architectures better fitted to problem. While less scaleable compared to the implementations with the previous hardware platforms, FPGAs excel when working with very limited resources, such as simulations running on an off-the-shelf personal computer.
@@ 10. Conclusions
Quantum computer simulators play an extremely important role in helping the development of new algorithms and hardware for the promising quantum computing paradigm. While several approaches and simulators have been proposed, the characteristics of the problem make it extremely hard to scale to systems with larger number of resources, required for solving more complex problems.

### v2

@@ 1. Introduction
While several simulators have recently been deployed and made publicly available, the need to simulate more complex algorithms, circuit configurations, or more accurately model qubit behavior leads to an exponential increase in computational and memory demand for the simulations. Scaling up the simulations without exponentially increasing the amount of necessary resources, requires optimisations or approximations at different levels. As such, several classical computer techniques, such as data compression or optimized parallel execution, have been applied to reduce the memory requirements and accelerate the computations (Lykov et al., 2022; Bayraktar et al., 2023).
@@
The organization of this survey is as follows. Section 2 introduces the essential concepts in quantum computing necessary for understanding the rest of the sections. In Section 3 we give an overview of the various parts of a quantum computer. Section 4 covers simulations at different levels and provides an overview of the available simulators for each of them, also introducing the main bottlenecks for scaling up the number of simulated qubits. Section 5 details the most common hardware platforms used for simulation. The following sections cover different acceleration methods, sorted by platform: CPU in Section 6, GPU in Section 7, and FPGA in Section 8. A summary and indication of future directions for hardware-aware optimization is presented in Section 9, and Section 10 concludes this work.
@@ 2.2.1. Single-qubit gates
These matrices correspond to the rotation of $\pi$ radians around respectively the x,y, and z axes of the Bloch sphere.
@@ 2.3. Quantum circuit execution
[Two qubit circuit, similar to two single-qubit circuits stacked vertically.]The image contains two quantum circuit diagrams, each depicting a different quantum gate operation applied to qubit states. The first diagram shows an initial quantum state psi on the left side of the X gate, represented as a rectangular box. The output of this operation is a transformed quantum state psi first on the right side of the box. The second diagram is similar, showing a quantum state labelled phi connected to the quantum gate Y also represented as a box. The output is a transformed quantum state labelled phi’.
@@ 3. Quantum computers
[Quantum computer control electronics(host, pulse generation, readout), interface blocks (DAC, ADC) and quantum chip]The image shows how the different parts of the system are connected. The different parts are placed in an elliptic configuration, starting from the host on the left and proceeding counter-clockwise back to the host. Starting from the host, represented as a computer symbol. An arrow goes from the host to the next component, the pulse generation block, which is represented as a rectangular box, situated on the bottom right of the host. The pulse generation block is then connected via an arrow to the base of the triangle used to represent the digital-to-analog converter. The point of the triangle, on the opposite side of the base, is on the right side, and it is connected with an arrow to the quantum chip, which is on the rightmost side of the figure. The quantum chip, represented as a chip symbol, is then connected to the analog-to-digital converter, which is represented as the same triangle as the DAC, but pointing in the opposite direction, and it is placed on the top left of the quantum chip. The ADC is then connected to the readout, represented as a rectangular box, and it is then connected back to the host.
@@ 3.1. Quantum chip
As already mentioned in the introduction, one of the main limitations of current physical qubits is their short coherence time. This is the timescale of the exponential decay of the qubit superposition state (Abad et al., 2022). The correctness of quantum computers is usually measured in terms of the fidelity of the operations or the output state.
@@ 4. Simulations of quantum computers
  • Algorithmic level — this is the highest level of the system, where algorithms are mapped to quantum circuits. This is comparable to the functional-level classical simulation.
@@ 4.2. Gate level
[The figure represents two coupled qubits, each depicted as a Josephson junction and a capacitance, and connected between them by a capacitor.]Each qubit is shown as a block with the labels omega1 or omega2 and alpha1 or alpha2. Each block has a Jhosepson junction on the left, represented as a square box with lines crossing from corner to corner, and a capacitor on the right, represented with the standard symbol. On the bottom the box is shown as connected to ground and on the top is connected to the other box through a capacitive coupling, shown as a capacitor labeled g12. The qubits are placed one on the left and one on the right, with the capacitive coupling in the top middle of the figure.
@@ 4.3. Algorithmic level
  • Intel Quantum SDK (Khalate et al., 2022): Provides two different simulators with two different modes of simulation, Generic Qubits and Intel Hardware. The Generic Qubits simulation is a state-vector-based simulation, which uses the Intel Quantum Simulator as a backend, allowing for a qubit-agnostic execution. The Intel Hardware mode uses as its core the Quantum Dot Simulator, which simulates the Intel quantum hardware, currently under development.
@@ 8. Acceleration using FPGA
Although FPGAs are more limited in terms of resources and frequency when compared to CPUs and GPUs, they allow for more flexibility, leading to architectures better fitted to problem. While less scalable compared to the implementations with the previous hardware platforms, FPGAs excel when working with very limited resources, such as simulations running on an off-the-shelf personal computer.
@@ 10. Conclusions
Quantum computer simulators play an extremely important role in helping the development of new algorithms and hardware for the promising quantum computing paradigm. While several approaches and simulators have been proposed, the characteristics of the problem make it extremely hard to scale to systems with a larger number of resources, required for solving more complex problems.
@@
Acknowledgements.
We acknowledge support from the Swedish Foundation for Strategic Research (grant number FUS21-0063), the Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA via the project 101113946 OpenSuperQPlus100, and from the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). AFK is also supported by the Swedish Research Council (grant number 2019-03696) and the Swedish Foundation for Strategic Research (grant number FFL21-0279). The work is also partially funded by the European High Performance Computing Joint Undertaking (JU) under Framework Partnership Agreement No 800928 and Specific Grant Agreement No 101036168 (EPI SGA2). The JU receives support from the European Union’s Horizon 2020 research and innovation programme and from Croatia, France, Germany, Greece, Italy, Netherlands, Portugal, Spain, Sweden, and Switzerland.